A cruise boat travels 48 miles downstream in 3 hours and returns upstream in 6 hours. Find the rate of the stream.
Answer by richwmiller(9144)

rate*time=distance
x*3=48
x=16 with current
y*6=48
y=6 against current
16-6=10 difference in speeds
10/2=5 mph stream current
5+11=16 with current
11-5=6 against current
11 mph with no current